OpenStudy (a.ahmed):

if 3^a = 4^b Find the value of this amount:- 9^a/b - 16^b/a

OpenStudy (anonymous):

@ShadowLegendX

OpenStudy (anonymous):

in such questions using LN helps so much. Are you ready to start thinking with me?

OpenStudy (a.ahmed):

ok i'm ready

OpenStudy (anonymous):

well, do you know what is LN? Also tell me what is the other form of the phrase below: \[\ln a^{n}=?\]

OpenStudy (a.ahmed):

i don't know what is LN

OpenStudy (unklerhaukus):

logarithm

OpenStudy (a.ahmed):

ok i know log but i don't know what is in ??

OpenStudy (a.ahmed):

what is the differance between log and in ?

OpenStudy (anonymous):

it is the Natural Logarithm. ok, do you know the Logarithm formulas? for example, can you write the phrase below in a simpler way: \[loga ^{N}=??\]

OpenStudy (a.ahmed):

i don't know can you help me

OpenStudy (anonymous):

OK, \[loga ^{N} = N.loga\] You got it?

OpenStudy (a.ahmed):

yes

OpenStudy (anonymous):

well! You have this equation: \[3^{a}=4^{b}\] Apply Log to both sides. What will be the result?

OpenStudy (anonymous):

by the sentence "Apply Log to both sides" I mean simply add the sign of logarithm to both sides in the equation above.

OpenStudy (a.ahmed):

ok . continue plz

OpenStudy (a.ahmed):

a. log 3 = b.log 4

OpenStudy (a.ahmed):

is that correct.?

OpenStudy (unklerhaukus):

yes, that's right, now solve for a/b

OpenStudy (a.ahmed):

i don't know how to solve for a/b

OpenStudy (unklerhaukus):

divide both sides by log 3, then divide both sides by b

OpenStudy (a.ahmed):

how??? can you sshow me how

OpenStudy (unklerhaukus):

a. log 3 = b.log 4 (dividing both sides of the equation by log 3) a. log 3 / log 3 = b.log 4 / log 3 (simplifying the lefthand side) a = b. log 4 / log 3 now divide by b

OpenStudy (unklerhaukus):

@A.ahmed

OpenStudy (a.ahmed):

great i'm sorry the connection went out

OpenStudy (a.ahmed):

you are amazing

OpenStudy (anonymous):

EXCUSE ME, A.Ahmed! Can you wait a minute so me and other viewers here talk about it?

OpenStudy (anonymous):

I guess there's a simpler way for solving it without using Logarithms. Please correct it if you see I'm making a mistake.

OpenStudy (zarkon):

like this... \[3^a = 4^b\] square both sides \[(3^a)^2=(4^b)^2\] \[3^{2a}=4^{2b}\] \[(3^2)^a=(4^2)^b\] \[(9)^a=(16)^b\] raise to the \(1/b\) power \[(9)^{a/b}=16\] raise to the \(1/a\) power \[9=(16)^{b/a}\] thus \[9^{a/b}-16^{b/a}=16-9=7\]

OpenStudy (anonymous):

\[3^{a}=4^{b} ==> 9^{a}=16^{b}\]

OpenStudy (anonymous):

YES! Exactly like what Zarkon said, without using Logarithm

OpenStudy (a.ahmed):

mr. yavar .i want to know the answer using logarithm .this answer is more diffucalr

OpenStudy (a.ahmed):

mr. @UnkleRhaukus can you continue the answer plz

OpenStudy (anonymous):

You're correct. Zarkon has given the simplest way to solve this. Using logarithm and its rules will make it more difficult to solve it.

OpenStudy (anonymous):

If you can understand it completely we can close this question, if you need complementary explanations on his solution, i can help you.

OpenStudy (a.ahmed):

mr. yavar can you continue the answer using logarithm by divide by b

OpenStudy (anonymous):

actually in the middle of solving it, the same solution came to my mind and i didn't continue it. Do you insist to know a solution using logarithms?

OpenStudy (a.ahmed):

yes

OpenStudy (anonymous):

Ok, so: \[a \log3 = b \log4\] SO \[\frac{ a }{ b }=\log_{3} 4\] AND\[\frac{ b }{ a }=\log_{4} 3\] now we can substitute these equations in the phrase below: 9^a/b - 16^b/a --> 3^2a/b-4^2b/a which after substitution becomes the attached file here. It doesn't show the equations well when I use it. So I have attached the file below:

OpenStudy (anonymous):

I have written the the most important formula I used, but I guess as long as you're not comfortable with logarithm rules, it's better to do it like Zarkon which is very much simpler. Do Great!

OpenStudy (a.ahmed):

you are great @Yavar . thank you too much

OpenStudy (anonymous):

:) you're welcome.